# Benchmarking pig detection and tracking under diverse and challenging conditions

### v1

@@ 5.1 Evaluated multi-object tracking models
We evaluated the tracking performance on the PigTrack test set using several widely adopted methods from the SORT family [81], ranging from classical approaches like ByteTrack [82] to more recent and stronger variants such as BoT-SORT [83]. These methods rely on bounding box outputs from a pre-trained detector that are linked across frames using motion or appearance-based association rules. In addition to these SORT-based models, we also benchmarked MOTRv2 [47] and MOTIP [48], two recent end-to-end trainable trackers, where detection and association are learned in a data-driven way. MOTRv2 and MOTIP achieve state-of-the-art results on the DanceTrack dataset [49], but have not yet been employed for pig tracking.
@@ 5.1.1 SORT family
In the original SORT model [81], detections are obtained independently for each frame of a video using Faster R-CNN [63]. The detections on the first two frames are matched with each other based on overlap to form object tracks. Then, each object’s motion is calculated as the spatial shift between the corresponding detections in the first two frames. Using these motion estimates, each object’s position in the third frame is predicted by applying the Kalman filter algorithm [84]. The actual detections on the third frame are then matched with existing tracks based on their IoU with the predicted object positions. This process is iterated for all frames to obtain tracks for the entire input video. The methods evaluated in this paper improve upon this basic paradigm in several ways.
@@
ByteTrack [82] addresses the fact that automatically discarding low-confidence detections leads to an increased number of false negatives. Therefore, the method uses a two-stage matching strategy. First, detections exceeding a predefined confidence threshold are associated with existing tracks as in the standard SORT paradigm. Then, detections falling below that threshold are associated with tracks that are still unmatched.
@@
OC-SORT [85] employs mechanisms to counteract noisy predictions that can arise when using the standard Kalman filter. For example, if an object has been occluded for several frames, even small errors in motion estimation can accumulate, leading to largely inaccurate predicted positions. OC-SORT tackles this by updating motion and position estimates based on the new observation once an object is matched again after a period of occlusion. This helps prevent the object from being lost in future frames due to inaccurate position estimates.
@@
Deep OC-SORT [86] builds upon OC-SORT and improves several key aspects. Most notably, it incorporates a deep association metric, yielding more robust tracking than motion-based association alone. Specifically, a pre-trained Re-ID model is applied on the detected objects to obtain appearance features. These features are then compared with those of existing tracks, and the resulting similarity scores are combined with motion information during the matching process.
@@
StrongSort [87] improves DeepSort [88], which was the first work to incorporate a deep association metric into SORT. Compared to its predecessor, StrongSort uses a more powerful object detector and Re-ID model. Other improvements include the calculation of an exponential moving average of the appearance features of a track. This makes the model less susceptible to noisy detections.
@@
BoT-SORT [83] builds upon the ByteTrack framework by incorporating a series of “bag of tricks” (hence BoT-SORT) that enhance both motion and appearance modeling. It employs a refined Kalman filter that results in more accurate position estimates. Furthermore, association is strengthened via an adaptive fusion of IoU and appearance similarity, allowing the tracker to dynamically weight motion and appearance cues based on whichever is more informative. BoT-SORT also introduces camera motion compensation to normalize global scene shifts.
@@
Improved Association (ImprAssoc [89]) also uses ByteTrack as a foundation and makes several improvements. Instead of using a two-stage matching strategy, all detections are matched in a single step with stricter similarity requirements for low-confidence detections. This allows the most fitting detections to be matched irrespective of their confidence score. Similar to BoT-Sort, ImprAssoc also takes both IoU and appearance similarity into account during matching. Lastly, detections with considerably high overlap are removed, as they likely represent duplicate predictions. This prevents the formation of ghost tracks that do not correspond to real objects.
@@ 5.1.2 End-to-end trainable models
Table 4: Model and training specifications for pig tracking. Pretrained Model indicates the dataset on which the model used for weight initialization was trained. The step learning rate schedules decay by a factor of 0.1 at each step. Training times for MOTIP full and MOTRv2 are reported for the default training setup that only uses the PigTrack training set (see Sec. 5.1.3). For training settings including more data, time scales roughly linearly with the number of frames.
Model Backbone Pretrained Model Batch Size Epochs LR Schedule LR (max/min) Train Time (h)
[…]
@@
MOTIP [48] uses a simple architecture built on top of DETR. Instead of using DETR to only output bounding boxes, it is configured to also output embedding vectors for each object. The embeddings of the first frame are concatenated with learned track ID embeddings and serve as context information for the tracking task. Object embeddings from the second frame retrieve this context information by interacting with the embeddings from the first frame via cross-attention. The enriched embeddings from the second frame are then used to predict the track ID using a linear layer. For later frames, MOTIP maintains a memory of object embeddings from multiple previous frames, enabling the model to recover identities even after temporary occlusions. Additionally, MOTIP reserves a special ID token that the model should predict whenever a previously unseen object enters the scene. The model can be trained by jointly optimizing the detection objective along with the ID prediction task.
MOTRv2 [47] is built on top of the MOTR framework [91], which also extends DETR for tracking. In MOTR, object embeddings are not only mapped to bounding boxes, but also serve as track queries that contain object-related information. These track queries interact with the image features of the next frame to predict new object embeddings and bounding boxes, enabling frame-by-frame tracking across a video. In addition, separately learned detect queries are used to identify newly appearing objects. The key contribution of MOTRv2 is the incorporation of bounding-box priors from a pre-trained object detector. By allowing track and detect queries to extract information from these priors, the burden of object detection is reduced, enabling MOTRv2 to focus more effectively on association modeling. This leads to a substantial performance increase compared to the MOTR baseline.
@@ 5.1.3 Implementation details
For MOTIP, we first pre-trained the DETR-based detection module on the PigDetect training and validation set and the PigTrack training set. The model achieving the best detection accuracy (see Sec. 5.2) on the PigTrack validation set was used as initialization for training the full model. MOTRv2 was initialized with a model pre-trained on DanceTrack [49]. Bounding box priors for MOTRv2 were obtained from Co-DINO trained on the PigDetect training set, as described in Sec. 4.1.3.
We evaluated three training configurations for both MOTIP and MOTRv2. In the default setup (1), the models were trained and validated according to the PigTrack dataset splits described in Sec. 3.3, and the best-performing model in terms of HOTA (see Sec. 5.2) on the validation set was used to obtain test performance. In a second setup (2), the validation set was added to the training data, with only two videos held out for validation. In a third setup (3), all test videos except one were also included in the training set, again retaining the same two videos for validation. By repeating this process for each of the 25 test videos, test performance can be evaluated in a leave-one-out fashion similar to cross-validation. In all three settings, we also included the PigDetect dataset during training by generating pseudo tracks from the images, following previous works [92, 93, 47]
@@
For the SORT-based models (see Sec. 5.1.1), we also used Co-DINO trained on the PigDetect training set (see Sec. 4.1.3) as an object detector. The hyperparameters of all SORT-based methods were tuned via Bayesian optimization [94] using Weights and Biases [95], with the objective of maximizing HOTA on the PigTrack validation set. The best hyperparameter configurations identified through this optimization procedure were used to obtain test results. All SORT-based models were run with code from the BoxMOT repository (version 11.0.6 [96]).
@@ 5.2 Evaluation metrics
For the evaluation of tracking, two aspects need to be considered: (1) The accuracy of object detections on individual video frames and (2) the accuracy of their associations across time. To assess these components, we employ the metrics introduced by Luiten et al. [97]. As for object detection evaluation (see Sec. 4.2), the first step to construct these metrics is to establish a matching between predicted and ground-truth bounding boxes and to enforce a minimum IoU of $\alpha$ for all pairs. For a given $\alpha$, the set of all matched pairs is referred to as true positives ($\text{TP}_{\alpha}$), the set of all unmatched predictions as false positives ($\text{FP}_{\alpha}$), and the set of all unmatched ground truths as false negatives ($\text{FN}_{\alpha}$). The detection accuracy is then simply defined as the proportion of $\text{TP}_{\alpha}$:
@@
$\text{FPA}_{\alpha}\text{(}c\text{)}$ =$=\{k\in\text{TP}_{\alpha}\mid\text{prID}(k)=\text{prID}(c)\land\text{gtID}(k)\neq\text{gtID}(c)\}$
@@
$\text{FNA}_{\alpha}\text{(}c\text{)}$ =$=\{k\in\text{TP}_{\alpha}\mid\text{prID}(k)\neq\text{prID}(c)\land\text{gtID}(k)=\text{gtID}(c)\}$
@@
For the sake of completeness, we also report MOTA [98], which is also a measure for both detection and tracking performance, but has been criticized for exaggerating the importance of detection [97]. Furthermore, we report the IDF1 score [99], which is an alternative measure for association performance. Finally, we also report the number of ID switches (IDSW), i.e. the number of times where an object is assigned a different ID than in the previous frame. Note that IDSW should not be seen as a measure for association performance, as it ignores the duration of switches and penalizes cases where the ID switches back to the correct one.
@@ 6.1 Object detection results on PigDetect
Figure 3: Visualization of challenging detection scenarios with model predictions. Shown are zoomed-in views of particularly difficult regions in two PigDetect test images. False positives are indicated by red dashed boxes, false negatives by red dotted boxes. All predicted bounding boxes with a score $\geq 0.2$ are visualized.
Object detection performance on the PigDetect test set was benchmarked using YOLOX and YOLOv8, two recent real-time models, and Co-DINO, a state-of-the-art model optimized for detection quality rather than inference speed (see Sec. 4.1). Results for the YOLOX and YOLOv8 models were obtained for various model sizes. As expected, YOLOv8 outperforms YOLOX, reflecting the performance improvements associated with more recent real-time object detection models (Tab. 5). The results further suggest that additional gains can be achieved by using performance-optimized models instead of real-time detectors. For example, an increase of roughly 2 % AP can be observed when using Co-DINO, which is currently state of the art on the COCO object detection dataset, compared to YOLOv8-X. However, the inference speed on Co-DINO is considerably slower, which makes its usage in real-time applications more difficult. In the end, practitioners must decide which trade-off in terms of performance and inference speed is desirable for the application at hand.
@@
YOLOX-S 78.9 97.4 31.8 8.9
YOLOX-M 80.9 97.5 36.9 25.3
YOLOX-L 82.5 98.5 43.8 54.1
YOLOX-X 82.7 98.5 49.6 99.0
YOLOv8-S 80.3 97.8 29.9 11.1
YOLOv8-M 82.5 98.4 32.0 25.9
YOLOv8-L 83.2 98.6 40.3 43.6
YOLOv8-X 83.6 98.6 40.1 68.2

### v2

@@ 5.1 Evaluated multi-object tracking models
We evaluated the tracking performance on the PigTrack test set using several widely adopted methods from the SORT family [bewley2016simple], ranging from classical approaches like ByteTrack [zhang2022bytetrack] to more recent and stronger variants such as BoT-SORT [aharon2022bot]. These methods rely on bounding box outputs from a pre-trained detector that are linked across frames using motion or appearance-based association rules. In addition to these SORT-based models, we also benchmarked MOTRv2 [zhang_motrv2_2023] and MOTIP [gao2025multiple], two recent end-to-end trainable trackers, where detection and association are learned in a data-driven way. MOTRv2 and MOTIP achieve state-of-the-art results on the DanceTrack dataset [sun2022dancetrack], but have not yet been employed for pig tracking.
@@ 5.1.1 SORT family
In the original SORT model [bewley2016simple], detections are obtained independently for each frame of a video using Faster R-CNN [ren2015faster]. The detections on the first two frames are matched with each other based on overlap to form object tracks. Then, each object’s motion is calculated as the spatial shift between the corresponding detections in the first two frames. Using these motion estimates, each object’s position in the third frame is predicted by applying the Kalman filter algorithm [kalman1960new]. The actual detections on the third frame are then matched with existing tracks based on their IoU with the predicted object positions. This process is iterated for all frames to obtain tracks for the entire input video. The methods evaluated in this paper improve upon this basic paradigm in several ways.
@@
ByteTrack [zhang2022bytetrack] addresses the fact that automatically discarding low-confidence detections leads to an increased number of false negatives. Therefore, the method uses a two-stage matching strategy. First, detections exceeding a predefined confidence threshold are associated with existing tracks as in the standard SORT paradigm. Then, detections falling below that threshold are associated with tracks that are still unmatched.
@@
OC-SORT [cao2023observation] employs mechanisms to counteract noisy predictions that can arise when using the standard Kalman filter. For example, if an object has been occluded for several frames, even small errors in motion estimation can accumulate, leading to largely inaccurate predicted positions. OC-SORT tackles this by updating motion and position estimates based on the new observation once an object is matched again after a period of occlusion. This helps prevent the object from being lost in future frames due to inaccurate position estimates.
@@
Deep OC-SORT [maggiolino2023deep] builds upon OC-SORT and improves several key aspects. Most notably, it incorporates a deep association metric, yielding more robust tracking than motion-based association alone. Specifically, a pre-trained Re-ID model is applied on the detected objects to obtain appearance features. These features are then compared with those of existing tracks, and the resulting similarity scores are combined with motion information during the matching process.
@@
StrongSort [du2023strongsort] improves DeepSort [wojke2017simple], which was the first work to incorporate a deep association metric into SORT. Compared to its predecessor, StrongSort uses a more powerful object detector and Re-ID model. Other improvements include the calculation of an exponential moving average of the appearance features of a track. This makes the model less susceptible to noisy detections.
@@
BoT-SORT [aharon2022bot] builds upon the ByteTrack framework by incorporating a series of “bag of tricks” (hence BoT-SORT) that enhance both motion and appearance modeling. It employs a refined Kalman filter that results in more accurate position estimates. Furthermore, association is strengthened via an adaptive fusion of IoU and appearance similarity, allowing the tracker to dynamically weight motion and appearance cues based on whichever is more informative. BoT-SORT also introduces camera motion compensation to normalize global scene shifts.
@@
Improved Association (ImprAssoc [stadler2023improved]) also uses ByteTrack as a foundation and makes several improvements. Instead of using a two-stage matching strategy, all detections are matched in a single step with stricter similarity requirements for low-confidence detections. This allows the most fitting detections to be matched irrespective of their confidence score. Similar to BoT-Sort, ImprAssoc also takes both IoU and appearance similarity into account during matching. Lastly, detections with considerably high overlap are removed, as they likely represent duplicate predictions. This prevents the formation of ghost tracks that do not correspond to real objects.
@@ 5.1.2 End-to-end trainable models
MOTIP [gao2025multiple] uses a simple architecture built on top of DETR. Instead of using DETR to only output bounding boxes, it is configured to also output embedding vectors for each object. The embeddings of the first frame are concatenated with learned track ID embeddings and serve as context information for the tracking task. Object embeddings from the second frame retrieve this context information by interacting with the embeddings from the first frame via cross-attention. The enriched embeddings from the second frame are then used to predict the track ID using a linear layer. For later frames, MOTIP maintains a memory of object embeddings from multiple previous frames, enabling the model to recover identities even after temporary occlusions. Additionally, MOTIP reserves a special ID token that the model should predict whenever a previously unseen object enters the scene. The model can be trained by jointly optimizing the detection objective along with the ID prediction task.
MOTRv2 [zhang_motrv2_2023] is built on top of the MOTR framework [zeng2022motr], which also extends DETR for tracking. In MOTR, object embeddings are not only mapped to bounding boxes, but also serve as track queries that contain object-related information. These track queries interact with the image features of the next frame to predict new object embeddings and bounding boxes, enabling frame-by-frame tracking across a video. In addition, separately learned detect queries are used to identify newly appearing objects. The key contribution of MOTRv2 is the incorporation of bounding-box priors from a pre-trained object detector. By allowing track and detect queries to extract information from these priors, the burden of object detection is reduced, enabling MOTRv2 to focus more effectively on association modeling. This leads to a substantial performance increase compared to the MOTR baseline.
@@ 5.1.3 Implementation details
Table 4: Model and training specifications for pig tracking. Pretrained Model indicates the dataset on which the model used for weight initialization was trained. The Step LR schedules reduce the learning rate by a factor of 0.1 at each step. LR BB scale denotes the multiplier applied to the backbone learning rate relative to the base learning rate given by LR (max/min). Time (h) denotes the training time. For MOTIP full and MOTRv2, training times are reported for the default setup that only uses the PigTrack training set (see Sec. 5.1.3). For training settings including more data, time scales roughly linearly with the number of frames.
Model Backbone Pretrained Model Batch Size Epochs LR Schedule LR (max/min) LR BB Scale Time (h)
MOTIP DETR ResNet50[he2016deep] COCO 32 50 Step (at 20 and 35) $2\times 10^{-4}\;\big/\;2\times 10^{-6}$ 0.1 5.3
MOTIP full ResNet50 PigDetect+PigTrack 8 18 Step (at 10 and 14) $1\times 10^{-4}\;\big/\;1\times 10^{-6}$ 0.1 17.0
MOTRv2 ResNet50 DanceTrack 8 20 Step (at 10) $2\times 10^{-4}\;\big/\;2\times 10^{-5}$ 0.1 11.5
For MOTIP, we first pre-trained the DETR-based detection module on the PigDetect training and validation set and the PigTrack training set. The model achieving the best detection accuracy (see Sec. 5.2) on the PigTrack validation set was used as initialization for training the full model. MOTRv2 was initialized with a model pre-trained on DanceTrack [sun2022dancetrack]. Bounding box priors for MOTRv2 were obtained from Co-DINO trained on the PigDetect training set, as described in Sec. 4.1.3.
We evaluated three training configurations for both MOTIP and MOTRv2. In the default setup (1), the models were trained and validated according to the PigTrack dataset splits described in Sec. 3.3, and the best-performing model in terms of HOTA (see Sec. 5.2) on the validation set was used to obtain test performance. In a second setup (2), the validation set was added to the training data, with only two videos held out for validation. In a third setup (3), all test videos except one were also included in the training set, again retaining the same two videos for validation. By repeating this process for each of the 25 test videos, test performance can be evaluated in a leave-one-out fashion similar to cross-validation. In all three settings, we also included the PigDetect dataset during training by generating pseudo tracks from the images, following previous works [zhou2020tracking, meinhardt2022trackformer, zhang_motrv2_2023]
@@
For the SORT-based models (see Sec. 5.1.1), we also used Co-DINO trained on the PigDetect training set (see Sec. 4.1.3) as an object detector, without additional fine-tuning on the PigTrack dataset. The hyperparameters of all SORT-based methods were tuned via Bayesian optimization [snoek2012practical] using Weights and Biases [wandb], with the objective of maximizing HOTA on the PigTrack validation set. The best hyperparameter configurations identified through this optimization procedure were used to obtain test results. All SORT-based models were run with code from the BoxMOT repository (version 11.0.6 [Brostrom_BoxMOT_pluggable_SOTA]).
@@ 5.2 Evaluation metrics
For the evaluation of tracking, two aspects need to be considered: (1) The accuracy of object detections on individual video frames and (2) the accuracy of their associations across time. To assess these components, we employ the metrics introduced by luiten2021hota. As for object detection evaluation (see Sec. 4.3), the first step to construct these metrics is to establish a matching between predicted and ground-truth bounding boxes and to enforce a minimum IoU of $\alpha$ for all pairs. For a given $\alpha$, the set of all matched pairs is referred to as true positives ($\text{TP}_{\alpha}$), the set of all unmatched predictions as false positives ($\text{FP}_{\alpha}$), and the set of all unmatched ground truths as false negatives ($\text{FN}_{\alpha}$). The detection accuracy is then simply defined as the proportion of $\text{TP}_{\alpha}$:
@@
$\text{FPA}_{\alpha}\text{(}c\text{)}$ $=\{k\in\text{TP}_{\alpha}\mid\text{prID}(k)=\text{prID}(c)\land\text{gtID}(k)\neq\text{gtID}(c)\}$
@@
$\text{FNA}_{\alpha}\text{(}c\text{)}$ $=\{k\in\text{TP}_{\alpha}\mid\text{prID}(k)\neq\text{prID}(c)\land\text{gtID}(k)=\text{gtID}(c)\}$
@@
For the sake of completeness, we also report MOTA [bernardin2008evaluating], which is also a measure for both detection and tracking performance, but has been criticized for exaggerating the importance of detection [luiten2021hota]. Furthermore, we report the IDF1 score [ristani2016performance], which is an alternative measure for association performance. We refer to the original publications for detailed definitions of MOTA and IDF1, as a full explanation of all metrics is beyond the scope of this work. Finally, we also report the number of ID switches (IDSW), i.e. the number of times where an object is assigned a different ID than in the previous frame. Note that IDSW should not be seen as a measure for association performance, as it ignores the duration of switches and penalizes cases where the ID switches back to the correct one.
@@ 6.1 Object detection results on PigDetect
Figure 3: Visualization of challenging detection scenarios with model predictions. Shown are zoomed-in views of particularly difficult regions in two PigDetect test images. False positives are indicated by red dashed boxes, false negatives by red dotted boxes. All predicted bounding boxes with a score $\geq 0.2$ are visualized, since low-confidence detections are frequently utilized in downstream tasks, for example in multi-object tracking [zhang2022bytetrack, aharon2022bot].
Object detection performance on the PigDetect test set was benchmarked using YOLOX, YOLOv8 and RT-DETR, three recent real-time models, and Co-DINO, a state-of-the-art model optimized for detection quality rather than inference speed (see Sec. 4.1). Results for YOLOX, YOLOv8 and RT-DETR were obtained for various model sizes. As expected, YOLOv8 outperforms YOLOX, reflecting the performance improvements associated with more recent real-time object detection models (Tab. 5). In contrast to the results reported by zhao2024detrs on the COCO dataset, RT-DETR performs worse than YOLOv8 on PigDetect and achieves results comparable to the older YOLOX model. Nonetheless, its speed-accuracy trade-off is more favorable. For example, RT-DETR-R50 matches YOLOX-L in accuracy while running twice as fast. The results further suggest that additional gains can be achieved by using state-of-the-art models instead of real-time detectors. For example, an increase of roughly 2 % AP can be observed when using Co-DINO, which is currently state of the art on the COCO object detection dataset, compared to YOLOv8-X. However, the inference speed of Co-DINO is considerably slower, which makes its usage in real-time applications more difficult. In the end, practitioners must decide which trade-off in terms of performance and inference speed is desirable for the application at hand. It should be noted that RT-DETR and Co-DINO might perform increasingly well with larger training corpora, as transformer-based architectures typically reach their full potential at scale.
@@
YOLOX-S 78.9 97.4 32 8.9
YOLOX-M 80.9 97.5 37 25.3
YOLOX-L 82.5 98.5 44 54.1
YOLOX-X 82.7 98.5 50 99.0
YOLOv8-S 80.3 97.8 30 11.1
YOLOv8-M 82.5 98.4 32 25.9
YOLOv8-L 83.2 98.6 40 43.6
YOLOv8-X 83.6 98.6 40 68.2
RT-DETR-R18 80.8 98.1 16 20.1
RT-DETR-R34 80.6 98.1 17 31.4
RT-DETR-R50 82.6 98.3 22 42.9
RT-DETR-R101 82.5 98.4 29 76.6
@@
6.2 Object detection results on third-party dataset
melfsen2023describing introduced a pig object detection dataset comprising images from a single pen environment captured from above at an angle of roughly $45^{\circ}$. The test partition of this dataset comprises 120 images, on which we evaluated the performance of the models trained on PigDetect (Tab. 6). All models successfully detect the majority of pigs ($\text{AP}_{0.5}$ ranging from 98.8 to 98.9). When taking into account localization accuracy, larger models generally achieve better performance, consistent with our observations on PigDetect. The best-performing models are RT-DETR-R50 and Co-DINO (AP = 78.6). Overall, performance differences between methods are less pronounced than on PigDetect, suggesting that the dataset contains less challenging examples to clearly differentiate between methods.
Interestingly, melfsen2023describing reported a considerably higher performance using YOLOv7 (AP = 86.4). Upon inspection of their training set, we found that this discrepancy is most likely due to near-identical training and test images (see Fig. 9), which make the results incomparable to ours. It appears that, to a large extent, the training and test sets were sampled from the same sequences without ensuring temporal disjointness.
Beyond this evaluation, we also applied Co-DINO to an unseen video dataset to obtain detections for tracking. This resulted in substantially improved tracking performance compared to the method proposed by the authors of the video dataset (see Sec. 6.5). Examples of predicted bounding boxes on this video dataset, the PigDetect test set, and the test set of melfsen2023describing are shown in Fig. 10. These qualitative results further corroborate the fact that models trained on PigDetect can provide high-quality detections in previously unseen environments.
Table 6: Performance measures for different detection methods on the test set published by melfsen2023describing. Best results are in bold, second-best underlined.
Model AP ↑ $\text{AP}_{0.5}$
Trained on data from [melfsen2023describing]
YOLOv7-default 86.4 -
Trained on PigDetect
YOLOX-S 76.2 98.9
YOLOX-M 77.2 98.9
YOLOX-L 77.7 98.9
YOLOX-X 77.7 98.9
YOLOv8-S 76.4 98.8
YOLOv8-M 77.6 98.9
YOLOv8-L 77.7 98.9
YOLOv8-X 77.6 98.9
RT-DETR-R18 76.6 98.9
RT-DETR-R34 77.5 98.9
RT-DETR-R50 78.6 98.9
RT-DETR-R101 77.5 98.8
Co-DINO 78.6 98.9